OpenStudy (anonymous):

at julio party,which was due to start at 8.00pm,3/5 of the guests arrived at least one hour late,1/8 of the guest arrived at least half an hour early,and 22 guests arrived between 7.30 and 9.00pm a)what fraction of guests arrived between 7.30 and 9.00pm? b)how many were there at the party altogether? c)how many guest arrived at least one hour late?

OpenStudy (anonymous):

JULIOS PARTYS RAGE

OpenStudy (shane_b):

Just logically break it down: 3/5 of the guests arrived at 9pm 1/8 of the guests arrived at 7:30 Part a) 3/5 + 1/8 guests arrived between 7:30 and 9pm Part b) 3/5x + 1/8x = 22 -> Solve for x Part c) Using part b's answer, what is 3/5x?

OpenStudy (anonymous):

a) 1-(3/5)-(1/8)=11/40 b) (11/40)x=22 x=80 c) (3/5)80=48

OpenStudy (shane_b):

@linh: Look at TheMind's answer which is correct. I screwed up part A...probably read it too fast.